# Self Forcing: Bridging the Train-Test Gap in Autoregressive Video Diffusion

### v1

@@ Abstract
We introduce Self Forcing, a novel training paradigm for autoregressive video diffusion models. It addresses the longstanding issue of exposure bias, where models trained on ground-truth context must generate sequences conditioned on their own imperfect outputs during inference. Unlike prior methods that denoise future frames based on ground-truth context frames, Self Forcing conditions each frame’s generation on previously self-generated outputs by performing autoregressive rollout with key-value (KV) caching during training. This strategy enables supervision through a holistic loss at the video level that directly evaluates the quality of the entire generated sequence, rather than relying solely on traditional frame-wise objectives. To ensure training efficiency, we employ a few-step diffusion model along with a stochastic gradient truncation strategy, effectively balancing computational cost and performance. We further introduce a rolling KV cache mechanism that enables efficient autoregressive video extrapolation. Extensive experiments demonstrate that our approach achieves real-time streaming video generation with sub-second latency on a single GPU, while matching or even surpassing the generation quality of significantly slower and non-causal diffusion models. footnotetext: Correspondence: xuhuang@adobe.com
@@ 3.2 Autoregressive Diffusion Post-Training via Self-Rollout
Nevertheless, implementing Self Forcing with standard many-step diffusion models would be computationally prohibitive, as it requires unrolling and backpropagation through long denoising chains. Therefore, we choose to use a few-step diffusion model $G_{\theta}$ to approximate each conditional distribution $p_{\theta}(x^{i}|x^{<i})$ in the autoregressive factorization. Consider $\{t_{0}=0,t_{1},\dots,t_{T}=1000\}$ a subsequence of timesteps $[0,...,1000]$, at each denoising step $t_{j}$ and frame index $i$, the model denoises an intermediate noisy frame $x^{i}_{t_{j}}$ conditioned on previous clean frames $x^{<i}$. It then injects Gaussian noise with a lower noise level into the denoised frame through the forward process $\Psi$ to obtain the noisy frame $x^{i}_{t_{j-1}}$ as the input to the next denoising step, following the standard practice in few-step diffusion models song2023consistency ; yin2024improved . The model distribution $p_{\theta}(x^{i}|x^{<i})$ is implicitly defined as $f_{\theta,t_{1}}\circ f_{\theta,t_{2}}\circ...\circ f_{\theta,t_{T}}(x^{i}_{t_{T}})$, where $f_{\theta,t_{j}}(x_{t_{j}}^{i})=\Psi(G_{\theta}(x^{i}_{t_{j}},t_{j},x^{<i}),\epsilon_{t_{j-1}},t_{j-1})$, and $x^{i}_{t_{T}}\sim\mathcal{N}(0,I)$.
@@
2:Number of video frames $N$
@@
15:                 $\mathbf{KV}.\mathrm{pop}(0)$triangleright Cache eviction
@@ 3.3 Holistic Distribution Matching Loss
  • Generative Adversarial Networks (GANs) goodfellow2014generative : It approximates the Jensen-Shannon divergence through a minimax game between the generator (our autoregressive diffusion model) and a discriminator that distinguishes between real and generated videos.
@@ 3.4 Long Video Generation with Rolling KV Cache
However, naive implementation of this mechanism leads to severe flickering artifacts due to distribution mismatch. Specifically, the first latent frame has different statistical properties than other frames: it only encodes the first image without performing temporal compression. The model, having always seen the first frame as the image latent during training, fails to generalize when the image latent is no longer visible in the rolling KV cache scenario. While we experimented with strategies similar to StreamingLLM streamingllm , keeping the first frame fixed while rolling KV cache of other frames, these proved ineffective for video generation. Our solution is straightforward but effective: during training, we restrict the attention window so the model cannot attend to the first chunk when denoising the final chunk, thereby simulating the conditions encountered during long video generation.
@@ Acknowledgments
We thank Tianwei Yin, Beidi Chen, Kaiwen Zheng, Kai Zhang, Gaurav Parmar, Yi Gu, Sai Bi, and Jianming Zhang for valuable discussions.
@@ Appendix A Implementation Details
We keep the preconditioning coefficients the same as the base models’ configuration, i.e., $c_{\text{skip}}=c_{\text{in}}=c_{\text{out}}=1$ and $c_{\text{noise}}(t)=t$. Our few-step diffusion process employs a uniform 4-step schedule $[t_{4},t_{3},t_{2},t_{1}]=[1000,750,500,250]$.
@@
[…]
@@
[…]
Generator/critic update ratio $5$ $5$ $1$
@@
where $x_{t}\sim p_{\text{data},t}$, $\hat{x}_{t}\sim p_{\theta,t}$ are the noisy real and fake data, respectively, $\epsilon$ and $\hat{\epsilon}$ are Gaussian noise sampled from $\mathcal{N}(0,1)$, and $f_{\psi}$ is the critic network (discriminator) of GAN. We use $\lambda=30$, $\sigma=0.05$ for all experiments. For a video generated from the output of the $s$-th step (see Algorithm 1 for details), we find that only sampling $t$ from $[t_{s-1},t_{s}]$ helps stabilize the training. We also adopt a large batch size of 768 for training stability.

### v2

@@ Abstract
We introduce Self Forcing, a novel training paradigm for autoregressive video diffusion models. It addresses the longstanding issue of exposure bias, where models trained on ground-truth context must generate sequences conditioned on their own imperfect outputs during inference. Unlike prior methods that denoise future frames based on ground-truth context frames, Self Forcing conditions each frame’s generation on previously self-generated outputs by performing autoregressive rollout with key-value (KV) caching during training. This strategy enables supervision through a holistic loss at the video level that directly evaluates the quality of the entire generated sequence, rather than relying solely on traditional frame-wise objectives. To ensure training efficiency, we employ a few-step diffusion model along with a stochastic gradient truncation strategy, effectively balancing computational cost and performance. We further introduce a rolling KV cache mechanism that enables efficient autoregressive video extrapolation. Extensive experiments demonstrate that our approach achieves real-time streaming video generation with sub-second latency on a single GPU, while matching or even surpassing the generation quality of significantly slower and non-causal diffusion models.
@@ 3.2 Autoregressive Diffusion Post-Training via Self-Rollout
Nevertheless, implementing Self Forcing with standard many-step diffusion models would be computationally prohibitive, as it requires unrolling and backpropagation through long denoising chains. Therefore, we choose to use a few-step diffusion model $G_{\theta}$ to approximate each conditional distribution $p_{\theta}(x^{i}|x^{<i})$ in the autoregressive factorization. Consider $\{t_{0}=0,t_{1},\dots,t_{T}=1000\}$ a subsequence of timesteps [0,,1000][0,...,1000], at each denoising step $t_{j}$ and frame index $i$, the model denoises an intermediate noisy frame $x^{i}_{t_{j}}$ conditioned on previous clean frames $x^{<i}$. It then injects Gaussian noise with a lower noise level into the denoised frame through the forward process $\Psi$ to obtain the noisy frame $x^{i}_{t_{j-1}}$ as the input to the next denoising step, following the standard practice in few-step diffusion models song2023consistency; yin2024improved. The model distribution $p_{\theta}(x^{i}|x^{<i})$ is implicitly defined as $f_{\theta,t_{1}}\circ f_{\theta,t_{2}}\circ...\circ f_{\theta,t_{T}}(x^{i}_{t_{T}})$, where $f_{\theta,t_{j}}(x_{t_{j}}^{i})=\Psi(G_{\theta}(x^{i}_{t_{j}},t_{j},x^{<i}),\epsilon_{t_{j-1}},t_{j-1})$, and $x^{i}_{t_{T}}\sim\mathcal{N}(0,I)$.
@@
2:Number of video frames NN
@@
15:      $\mathbf{KV}.\mathrm{pop}(0)$ $\triangleright$ Cache eviction
@@ 3.3 Holistic Distribution Matching Loss
  • Generative Adversarial Networks (GANs) goodfellow2014generative: It approximately minimizes the Jensen-Shannon divergence through a minimax game between the generator (our autoregressive diffusion model) and a discriminator that distinguishes between real and generated videos.
@@ 3.4 Long Video Generation with Rolling KV Cache
However, naive implementation of this mechanism leads to severe flickering artifacts due to distribution mismatch. Specifically, the first latent frame has different statistical properties than other frames: it only encodes the first image without performing temporal compression. The model, having always seen the first frame as the image latent during training, fails to generalize when the image latent is no longer visible in the rolling KV cache scenario. Our solution is straightforward but effective: during training, we restrict the attention window so the model cannot attend to the first chunk when denoising the final chunk, thereby simulating the conditions encountered during long video generation.
@@ Acknowledgments
We thank Tianwei Yin, Beidi Chen, Kaiwen Zheng, Kai Zhang, Gaurav Parmar, Yi Gu, Sai Bi, and Jianming Zhang for valuable discussions. G. He and M. Zhou acknowledge the support of NSF-IIS 2212418 and NIH-R37 CA271186.
@@ Appendix A Implementation Details
We keep the preconditioning coefficients the same as the base models’ configuration, i.e., $c_{\text{skip}}=c_{\text{in}}=c_{\text{out}}=1$ and $c_{\text{noise}}(t)=t$. Our few-step diffusion process employs a uniform 4-step schedule $t_{4},t_{3},t_{2},t_{1}]=[1000,750,500,250]$.
@@
$\epsilon=1\text{e-8}$, weight_decay=0.01=0.01 $\epsilon=1\text{e-8}$, weight_decay=0=0 $\epsilon=1\text{e-8}$, weight_decay=0.01=0.01
@@
$\epsilon=1\text{e-8}$, weight_decay=0.01=0.01 $\epsilon=1\text{e-8}$, weight_decay=0=0 $\epsilon=1\text{e-8}$, weight_decay=0.01=0.01
Learning rate ($G_{\theta}$) $2\text{e-6}$ $2\text{e-6}$ $2\text{e-6}$
Learning rate ($f_{\psi}$) $4\text{e-7}$ $2\text{e-6}$ $2\text{e-6}$
Generator/critic update ratio 55 55 11
@@
where $x_{t}\sim p_{\text{data},t}$, $\hat{x}_{t}\sim p_{\theta,t}$ are the noisy real and fake data, respectively, $\epsilon$ and $\hat{\epsilon}$ are Gaussian noise sampled from $\mathcal{N}(0,1)$, and $f_{\psi}$ is the critic network (discriminator) of GAN. We use $\lambda=30$, $\sigma=0.05$ for all experiments. For a video generated from the output of the $s$-th step (see Algorithm 1 for details), we find that only sampling $t$ from $t_{s-1},t_{s}]$ helps stabilize the training. We also adopt a large batch size of 768 for training stability.